A 8.10 kg rock whose density is 4100 kg/m3 is suspended by a string such that half of the rock's volume is under water.

What is the tension in the string?
Answer #1

Given,

m = 8.1 kg ; rho = 4100 kg/m^3

V = m/d = 8.1/4100 = 1.98 x 10^-3 m^3

the forces on rock are:

W = mg

W = 8.1 x 9.8 = 79.38 N

Fb = V rho g

Volume of dispalced water:

V' = V/2 = 1.98 x 10^-3/2 = 9.9 x 10^-4 m^3

Fb = 9.9 x 10^-4 x 1000 x 9.8 = 9.68 N

Tenstion T which acts upwards,under equilibrium

Fnet = 0

Fb + T - W = 0

T = W - Fb

T = 79.38 - 9.68 = 69.7 N

Hence, T = 69.7 N